OpenStudy (anonymous):

Your piggy bank contains $19.75 in dimes and quarters. There are 100 coins in all. How many dimes are there?

OpenStudy (cj49):

Let x = be the number of dimes and y = number of quarters. Since the total number of coins is 100, x + y = 100. Since the total is cent value is 19.75, comprised of .10 for dimes and .25 for quarters, we have .10x + .25y = 19.75. Using substitution: x = 100 - y Plug this into the other equation 10 (100 - y) + .25 y = 19.75 10 - .10 y + .25 y = 19.75 10 + .15y = 19.75 .15 y = 9.75 y = 65 So there are 65 quarters. Since x + y = 100, there must be 35 dimes.

OpenStudy (anonymous):

I'm sorry but i got confused when substituting x = 100 - y